OpenStudy (anonymous):

What is the mass of 3.5 moles of NaCl? A. 5.8 × 101 grams B. 3.5 × 100 grams C. 2.0 × 102 grams D. 6.2 × 101 grams E. 6.0 × 10–2 grams

OpenStudy (chmvijay):

which one do u thin k

OpenStudy (anonymous):

I was thinking A?

OpenStudy (jfraser):

what does 1 mole of NaCl weigh? the periodic table should be your best friend right about now

OpenStudy (anonymous):

Na = 22.99 g/mol

thomaster (thomaster):

yep and Cl = 35.453 g/mol so NaCl = 22.989 + 35.453 = 58.442 g/mol now multiply that by the number of moles

OpenStudy (anonymous):

Is C correct?

thomaster (thomaster):

58.442 * 3.5 = 204.547 grams If you write that in scientific notation you'll get 2.04 x 10^2 grams So C is correct

OpenStudy (anonymous):

Thanks :)!